# Uniformly perfect measures on strictly convex planar graphs are $L^{2}$-flattening

### v1

@@ 4.1. Energy flattening
By the Fourier analytic expression for Riesz energy [22, equation (4.5)], Plancherel, and a dyadic frequency decomposition, our assumption implies that
$I_{\alpha+\eta}^{\delta}\left(\Pi^{2^{k_{0}}}\right)\leq\delta^{-\kappa}.$
This is our claim. ∎
The remaining task is to prove (4.3) with $k_{0}=\lceil 20/\kappa^{\prime}\rceil$. Before doing this, let us record that the sequence $\{J_{r}(k)\}_{k\in\mathbb{N}}$ is non-increasing. In fact, by Young’s inequality (or Plancherel), writing $\left\lVert\nu\right\rVert_{1}$ for both total variance and $L^{1}$-norm,
$J_{r}(k+1)=\left\lVert\Pi_{r}^{2^{k}}\ast\Pi_{r}^{2^{k}}\right\rVert_{2}\leq\left\lVert\Pi_{r}^{2^{k}}\right\rVert_{1}\cdot\left\lVert\Pi_{r}^{2^{k}}\right\rVert_{2}=\left\lVert\Pi_{r}^{2^{k}}\right\rVert_{2}=J_{r}(k),\quad k\geq 0.$
Thus, it suffices to prove (4.3) for each $r\in[\delta,1]$ fixed, and for $k=k(r)\leq k_{0}$.
We first dispose of the case $r\geq\delta^{\frac{\kappa}{8}}$. Since $\alpha+\eta\leq 2$,
$J_{r}(0)\leq Cr^{-1}\leq Cr^{-2}\leq\delta^{-\frac{\kappa}{4}}r^{\frac{\alpha+\eta-2}{2}}.$
Thus, (4.3) is satisfied with $k=1$ for $r\geq\delta^{\kappa/8}$.
From now on we fix $r\in[\delta,\delta^{\frac{\kappa}{8}}]$.
@@
There either exists $k=k(r)\leq\lceil\frac{20}{\kappa^{\prime}}\rceil$ such that
$r^{\kappa^{\prime}/10}J_{r}(k)\leq J_{r}(k+1)\leq J_{r}(k),$ (4.4)
or otherwise (4.3) holds with $k=\lceil\tfrac{20}{\kappa^{\prime}}\rceil$ (provided $\delta>0$ is sufficiently small in terms of $R$).
@@
Suppose that for every $k\leq\lceil\frac{20}{\kappa^{\prime}}\rceil$ we have
$r^{\kappa^{\prime}/10}J_{r}(k)\leq J_{r}(k+1).$
Applying this $\lceil\frac{20}{\kappa^{\prime}}\rceil$ times we then have
$J_{r}(\lceil\tfrac{20}{\kappa^{\prime}}\rceil)\leq r^{(20/\kappa^{\prime})\cdot(\kappa^{\prime}/10)}J_{r}(0)\lesssim_{R}r^{2}\cdot r^{-2}=1.$
This implies (4.3) for $\delta=\delta(R)>0$ small enough. ∎
For the remainder of the proof, we may assume that the first option in Claim 4.4 holds: thus, there exists $k=k(r)\leq\lceil\tfrac{20}{\kappa^{\prime}}\rceil$ satisfying (4.4). We claim that (4.3) is satisfied with this choice of $k$.
We start by performing a discretisation at scale $r$ of the function $\Pi_{r}^{2^{k}}$. First, define
$a_{Q}:=\sup_{x\in Q}\Pi_{r}^{2k}(x),\qquad Q\in\mathcal{D}_{r}(\mathbb{R}^{2}).$
Now, define
$A_{0}:=\bigcup\{Q\in\mathcal{D}_{r}(\mathbb{R}^{2}):\,a_{Q}\leq 1\},$
and for $j\geq 1$,
$A_{0}:=\bigcup\{Q\in\mathcal{D}_{r}(\mathbb{R}^{2}):\,2^{j-1}\leq a_{Q}\leq 2^{j}\}.$
Note that the sets $A_{j}$ are disjoint for distinct $j$, and that $A_{j}=\emptyset$ for all $j\geq 4\log(\frac{1}{r})+1$ since $\left\lVert\Pi_{r}^{2^{k}}\right\rVert<r^{-2}$ for $k\geq 1$.
We require the following claim, which is [22, Claim 4.13]:
Claim 4.5.
There exists $j\geq 0$ and a set $A:=A_{j}$ such that:
  1. (1)
    We have $\left\lVert\Pi_{r}^{2^{k}}\right\rVert_{2}\lesssim|A|_{r}^{-1/2}r^{-1-\kappa^{\prime}/5}\leq|A|_{r}^{-1/2}r^{-1-\kappa/5}$;
  2. (2)
    $\Pi_{8r}^{2^{k}}(A)\gtrsim r^{\kappa^{\prime}/5}.$
By Claim 4.5 Part (2) and since $\operatorname{spt}\psi_{r}\subset B(r)$,
$\Pi^{2^{k}}\left([A]_{8r}\right)\geq\Pi_{8r}^{2^{k}}(A)\gtrsim r^{\kappa^{\prime}/5}.$ (4.5)
@@
$(\nu*\sigma)([A]_{8r}+z)\gtrsim r^{\kappa^{\prime}/5}/R^{2},\qquad z\in\mathbb{R}^{2}.$
So, recalling from (4.2) that $\kappa^{\prime}\leq\eta$, and taking $\delta>0$ small enough in terms of $R$,
$(\nu*\sigma)([A]_{8r}+z)\geq r^{\eta}.$
We also note that, using the hypothesis of the proposition, $r\leq\delta^{\kappa/8}$, and $\epsilon\leq\kappa\eta/8$,
$I_{\alpha}^{r}(\nu)\leq I_{\alpha}^{\delta}(\mu)\leq\delta^{-\epsilon}\leq r^{-\eta}.$
We may therefore apply Corollary 4.1 to the measures $\nu,\sigma$ and the set $E:=[A]_{8r}+z$, and conclude that
$|A|_{r}\geq r^{-\alpha-\eta}.$
So, by Claim 4.5,
$J_{r}(k)=\left\lVert\Pi_{r}^{2^{k}}\right\rVert_{2}\lesssim|A|_{r}^{-\frac{1}{2}}r^{-1-\kappa/5}\leq r^{\frac{\alpha+\eta}{2}}\cdot r^{-1-\kappa/5}\leq\delta^{-\kappa/5}r^{\frac{\alpha+\eta-2}{2}}.$
This completes the proof. ∎
The following corollary is a counterpart of [22, Corollary 4.18] in our setting:
@@
For all $D\geq 1$, $\mathfrak{d}>0$, $\beta\in[0,1)$, $t\in[0,2)$, and $\kappa>0$, there exist $k_{0}=k_{0}(D,\beta,\kappa,t)\in\mathbb{N}$ and $\delta_{0}=\delta_{0}(D,\mathfrak{d},\beta,\kappa,t)>0$ such that the following holds for all $\delta\in(0,\delta_{0}]$. Assume that $\sigma$ is a $(D,\beta)$-uniformly perfect probability measure supported on $\mathbb{P}$ with $\operatorname{diam}(\operatorname{spt}\sigma)\geq\mathfrak{d}$. Then,
$I^{\delta}_{t}(\sigma^{k})\leq\delta^{-\kappa},\qquad k\geq k_{0}.$
@@
Let $\alpha_{0}:=\tfrac{1}{2}\alpha$. Now the Frostman property easily implies (see [18, Chapter 8]) that
$I^{\delta}_{\alpha_{0}}(\sigma)\lesssim_{D,\mathfrak{d},\beta}1,\qquad\delta\in(0,1].$ (4.6)
We now proceed to define a sequence of exponents $\{\alpha_{j}\}_{j=0}^{\infty}$, where $\alpha_{0}>0$ is the exponent defined just above. Given $j\geq 0$, we then inductively define
@@
where $\eta_{j}:=\eta(D,\alpha_{j},\beta)>0$ is the constant provided by Proposition 4.2. Note that $\alpha_{j}\nearrow 2$ as $j\to\infty$, since the constants $\eta(D,\alpha,\beta)>0$ stay bounded away from zero as $\alpha$ ranges on any fixed compact subset of $[0,2)$. In particular, given $t\in[0,2)$ as in the statement of the corollary, there exists $j_{0}=j_{0}(D,\beta,t)\in\mathbb{N}$ such that $\alpha_{j_{0}}\geq t$.
We make the following claim, to be proved by induction. Fix $j\geq 0$ and $\kappa_{j}>0$. Then, there exist $k^{j}=k^{j}(D,\beta,j,\kappa_{j})\in\mathbb{N}$ and $\delta^{j}=\delta^{j}(D,\mathfrak{d},\beta,j,\kappa_{j})>0$ such that
$I_{\alpha_{j}}^{\delta}(\sigma^{k^{j}})\leq\delta^{-\kappa_{j}},\qquad\delta\in(0,\delta^{j}].$ (4.7)
Once this has been established, we may complete the proof of Corollary 4.6 by applying the inductive claim with $j:=j_{0}$ and $\kappa_{j_{0}}:=\kappa$. We infer the existence of an integer $k_{0}:=k^{j_{0}}\in\mathbb{N}$ and a scale $\delta_{0}:=\delta^{j_{0}}$ such that (4.7) holds for all $k\geq k_{0}$. Thus,
$I_{t}^{\delta}(\sigma^{k})\leq I_{\alpha_{j_{0}}}^{\delta}(\sigma^{k^{j_{0}}})\leq\delta^{-\kappa},\qquad k\geq k_{0},\,\delta\in(0,\delta_{0}],$
as claimed in the corollary.
Let us then prove the inductive claim. The case $j=0$ follows from (4.6), with $k^{0}\equiv 1$. Let us then assume that the claim has already been established for some $j\geq 0$. Fix $\kappa_{j+1}>0$. Apply Proposition 4.2 with the parameters $D,\beta$, $\alpha:=\alpha_{j}$, $\kappa:=\kappa_{j+1}$, and a radius $R_{j}=R(D,\alpha_{j},\beta,j,\kappa_{j+1})\geq 1$, which is complicated to define, but whose choice we nonetheless discuss right away to avoid suspicions of circular reasoning. Note that the constant
$\epsilon=\epsilon(\alpha_{j},\beta,\kappa_{j+1},D)>0$ (4.8)
in Proposition 4.2 is independent of "RR" (it is explicitly given in (4.2)). Therefore, we may use $\epsilon$ to define $R$. We do this by applying the inductive hypothesis (4.7) with parameter $\kappa_{j}:=\epsilon$. The conclusion is the existence of
$k^{j}=k^{j}(D,\beta,j,\epsilon)\in\mathbb{N}\quad\text{and}\quad\delta^{j}=\delta^{j}(D,\mathfrak{d},\beta,j,\epsilon)>0$
such that
$I_{\alpha_{j}}^{\delta}(\sigma^{k^{j}})\leq\delta^{-\epsilon},\qquad\delta\in(0,\delta^{j}].$ (4.9)
Since $\epsilon=\epsilon(D,\alpha_{j},\beta,\kappa_{j+1})>0$, the integer $k^{j}$ only depends on $D,\alpha_{j},\beta,j,\kappa_{j+1}$. Moreover, since $\operatorname{spt}\sigma\subset[-2,2]^{2}$, we have
$\operatorname{spt}\sigma^{k^{j}}\subset[-2k^{j},2k^{j}]^{2}.$ (4.10)
We now set $R_{j}:=2k^{j}$, which ensures that $\operatorname{spt}\sigma^{k_{j}}\subset[-R_{j},R_{j}]^{2}$.
Now that all the parameters $D,\beta,\alpha=\alpha_{j},\kappa=\kappa_{j+1}$, $R_{j}$ have been defined, we return to the application of Proposition 4.2. The conclusion is the existence of $\epsilon=\epsilon(\alpha_{j},\beta,\kappa_{j+1},D)>0$ (the constant we already encountered at (4.8)), $k_{0}=k_{0}(\alpha_{j},\beta,\kappa_{j+1},D)$, and $\delta_{0}:=\delta_{0}(\alpha_{j},\beta,j,\kappa_{j+1},\mathfrak{d},D,R_{j})>0$ such that the following holds for $\delta\in(0,\delta_{0}]$. If $\mu$ is a Borel probability measure satisfying
$\operatorname{spt}\mu\subset[-R_{j},R_{j}]^{2}\quad\text{and}\quad I_{\alpha_{j}}^{\delta}(\mu)\leq\delta^{-\epsilon},$ (4.11)
@@
$I^{\delta}_{\alpha_{j+1}}((\mu\ast\sigma)^{k_{0}})=I^{\delta}_{\alpha_{j}+\eta}((\mu\ast\sigma)^{k_{0}})\leq\delta^{-\kappa_{j+1}}.$ (4.12)
But now a combination of (4.9)-(4.10) shows that $\mu=\sigma^{k_{j}}$ satisfies (4.11) for $\delta\in(0,\delta^{j}]$, so (4.12) yields
$I_{\alpha_{j+1}}^{\delta}(\sigma^{k_{0}(k_{j}+1)})\leq\delta^{-\kappa_{j+1}},\qquad\delta\in(0,\min\{\delta_{0},\delta^{j}\}].$
@@
Starting from Corollary 4.6, the proof is similar to the proof of [22, Theorem 1.1] presented in [22, Section 5]. We thus leave out some computations.
@@
Fix $R\geq 1$, and let $\{\varphi_{\delta}\}_{\delta>0}$ be an approximate identity with the property $|\widehat{\varphi_{\delta}}(\xi)|\gtrsim 1$ for $|\xi|\leq\delta$. Then, writing $\delta:=R^{-1}$, and for $p=2k\in 2\mathbb{N}$, one may check that
$\|\hat{\sigma}\|_{L^{p}(B(R))}^{p}\lesssim_{u}R^{2-u}I^{\delta}_{u}(\sigma^{k}),\qquad u\in(0,2),$
where $I_{u}^{\delta}$ refers to the $\delta$-mollified energy defined relative to $\{\varphi_{\delta}\}_{\delta>0}$.
In particular, this holds for $u:=2-\epsilon/2$, where $\epsilon\in(0,1)$ is the parameter given in the statement. Write $\kappa:=\epsilon/2$. Then, according to Corollary 4.6, we have $I^{\delta}_{u}(\sigma^{k})\leq\delta^{-\kappa}=R^{\epsilon/2}$, provided $k\geq k_{0}(D,\beta,\epsilon)$, and $0<\delta\leq\delta_{0}(D,\mathfrak{d},\beta,\epsilon)$. For such $k$, and $p=2k$, we have now established
$\|\hat{\sigma}\|_{L^{p}(B(R))}^{p}\lesssim_{t}R^{2-u}I_{u}^{\delta}(\sigma^{k})\leq R^{\epsilon},\qquad R\geq 1/\delta_{0}.$
This completes the proof of Theorem 4.7. ∎
4.3. Proof of Corollary 1.3
Let us first recall the statement of Corollary 1.3; then, we give its (standard) proof.
Corollary 4.8.
For every $D>1$, $\beta\in[0,1)$, $\mathfrak{d}>0$, and $\epsilon\in(0,1)$ there exists $\kappa=\kappa(D,\beta,\epsilon)>0$ such that the following holds.
Let $R\geq 1$, $f\in L^{2}(\mathbb{R}^{2})$, and $\sigma$ a Borel probability measure such that:
  1. (1)
    $\operatorname{spt}\hat{f}\subset B(R)$.
  2. (2)
    $\sigma$ is a $(D,\beta)$-uniformly perfect measure on $\mathbb{P}$ such that $\operatorname{diam}(\operatorname{spt}\sigma)\geq\mathfrak{d}$.
Then,
$\|f\|_{L^{2}}\geq R^{\epsilon}\|f\|_{L^{1}}\quad\Longrightarrow\quad\|f\ast\sigma\|_{L^{2}}\lesssim_{D,\mathfrak{d},\beta,\eta}R^{-\kappa}\|f\|_{L^{2}}.$
Proof.
Apply Theorem 1.2 with constants $D,\beta,\epsilon$ to produce the constant $p=p(D,\beta,\epsilon)\geq 2$ such that (1.3) holds. Since $\sigma$ is a probability measure we have
$\left\lVert\hat{\sigma}\right\rVert_{L^{2}(B(R))}^{2}\leq\left\lVert\hat{\sigma}\right\rVert_{L^{1}(B(R))},$
whence we may assume $p\geq 2$. Thus, writing $q:=\frac{p}{2}$ and $q^{\prime}:=\frac{p}{p-2}$, we have:
$\|f\ast\sigma\|_{L^{2}}^{2}$ $=$ $\int_{B(R)}|\hat{f}\hat{\sigma}|^{2}$
$\leq$ $\Big{(}\int|\hat{f}|^{2q^{\prime}}\Big{)}^{1/q^{\prime}}\|\hat{\sigma}\|_{L^{p}(B(R))}^{2}$
$\leq$ $\|f\|_{L^{1}}^{2-2/q^{\prime}}\|f\|_{L^{2}}^{2/q^{\prime}}\cdot R^{2\epsilon/p}$
$=$ $\|f\|_{L^{1}}^{4/p}\|f\|_{L^{2}}^{2-4/p}\cdot R^{2\epsilon/p}$
$\leq$ $\|f\|_{L^{2}}^{2}\cdot R^{-2\epsilon/p}.$
The first equality follows from (1) and Plancherel, the second inequality is Hölder’s inequality with exponents $q$ and dual exponent $q^{\prime}$, the third inequality follows from Theorem 1.2 and yet another application of Hölder’s inequality, the fourth equality is just our choice of parameters, and the final inequlaity follows from our assumption on $\|f\|_{L^{2}}$.
We conclude that
$\|f\ast\sigma\|_{L^{2}}\leq R^{-\kappa}\|f\|_{L^{2}},\text{ with }\kappa:=\epsilon/p.$
The proof is complete. ∎

### v2

@@ 4.1. $L^{2}$-flattening
Suppose that $\delta^{\eta/10}J(k)\leq J(k+1)$ for every $k\leq\lceil 20/\eta\rceil$. Applying this $\lceil 20/\eta\rceil$ times,
$J(\lceil 20/\eta\rceil)\leq\delta^{(20/\eta)\cdot(\eta/10)}J(0)\lesssim_{R}\delta^{2}\cdot\delta^{-2}=1.$
This implies (4.3) for $\delta=\delta(R)>0$ small enough. ∎
For the remainder of the proof, we may assume that the first option in Claim 4.3 holds: there exists $k\leq\lceil 20/\eta\rceil$ satisfying (4.4). We claim that (4.3) is satisfied with this $k$.
We start by performing a discretisation at scale $\delta$ of the function $\Pi^{2^{k}}$. First, define
$a_{Q}:=\sup_{x\in Q}\Pi^{2k}(x),\qquad Q\in\mathcal{D}_{\delta}(\mathbb{R}^{2}).$
Now, define
$A_{0}:=\bigcup\{Q\in\mathcal{D}_{\delta}(\mathbb{R}^{2}):\,a_{Q}\leq 1\},$
and for $j\geq 1$,
$A_{0}:=\bigcup_{j}\{Q\in\mathcal{D}_{\delta}(\mathbb{R}^{2}):\,2^{j-1}\leq a_{Q}\leq 2^{j}\}.=:\bigcup_{j}A_{j}.$ (4.5)
Note that the sets $A_{j}$ are disjoint for distinct $j$, and $A_{j}=\emptyset$ for all $j\geq C\log(\frac{1}{\delta})$ (for $C\geq 1$ absolute) since $\big\|\Pi^{2^{k}}\big\|\lesssim\delta^{-2}$ for $k\geq 1$.
The following claim is [22, Claim 4.13], but we repeat the details for completeness.
@@
There exists $j\in\{0,\ldots,C\log(1/\delta)\}$ such that writing $A:=A_{j}$:
  1. (1)
    $\big\|\Pi^{2^{k}}\big\|_{2}\lesssim_{\eta}|A|_{\delta}^{-1/2}\delta^{-1-\eta/5}$;
  2. (2)
    $\Pi_{8\delta}^{2^{k}}(A)\gtrsim_{\eta}\delta^{\eta/5}.$
@@
The next estimate follows from (4.5), and $a_{Q}\lesssim\inf_{x\in Q}\Pi_{8\delta}^{2^{k}}(x)$:
$\Pi^{2^{k}}\leq\sum_{j=0}^{C\log(1/\delta)}2^{j}\cdot\mathbf{1}_{A_{j}}\quad\text{and}\quad 2^{j}\cdot\mathbf{1}_{A_{j}}\lesssim C\Pi_{8\delta}^{2^{k}}\text{ for all }j\geq 0.$ (4.6)
(See [22, (4.11)-(4.12)] for full details.) Taking $L^{2}$-norms, using $\psi_{r}\lesssim\psi_{r}\ast\psi_{r}$, and applying the triangle inequality, we may pigeonhole an index $j\geq 0$ and a set $A=A_{j}$ such that
$\|\Pi^{2^{k+1}}\|_{2}\lesssim\|\Pi^{2^{k}}*\Pi^{2^{k}}\|_{2}\lesssim\log(1/\delta)\cdot 2^{j}\|\mathbf{1}_{A}*\Pi^{2^{k}}\|_{2}.$
On the other hand, by Plancherel and Cauchy-Schwarz,
$\|\mathbf{1}_{A}*\Pi_{r}^{2^{k}}\|_{2}\leq\|\mathbf{1}_{A}*\mathbf{1}_{A}\|_{2}^{1/2}\|\Pi^{2^{k+1}}\|_{2}^{1/2}.$
Combining these two estimates with (4.4) yields
$\delta^{\eta/10}\|\Pi^{2^{k}}\|_{2}\leq\|\Pi^{2^{k+1}}\|_{2}\lesssim(\log(1/\delta))^{2}2^{2j}\|\mathbf{1}_{A}*\mathbf{1}_{A}\|_{2}\lesssim_{\eta}\delta^{-\eta/10}2^{2j}\|\mathbf{1}_{A}\|_{1}\|\mathbf{1}_{A}\|_{2}.$ (4.7)
Here, $A$ is a union of elements in $\mathcal{D}_{\delta}(\mathbb{R}^{2})$, so $\|\mathbf{1}_{A}\|_{1}=\delta^{2}|A|_{\delta}$ and $\|\mathbf{1}_{A}\|_{2}=\delta|A|_{\delta}^{1/2}$. In particular, combining (4.7) with $2^{j}\|\mathbf{1}_{A}\|_{1}\lesssim\|\Pi_{8\delta}^{2^{k}}\|_{1}\leq 1$ (see (4.6)), we obtain
$\|\Pi^{2^{k}}\|_{2}\lesssim_{\eta}\delta^{-\eta/5}2^{2j}\|\mathbf{1}_{A}\|_{1}\|\mathbf{1}_{A}\|_{2}\lesssim\delta^{-\eta/5}\|\mathbf{1}_{A}\|_{1}^{-1}\|\mathbf{1}_{A}\|_{2}\leq\delta^{-\eta/5-1}|A|_{\delta}^{-\frac{1}{2}},$ (4.8)
concluding the proof of Claim 4.4(1).
Moving towards proving part (2), we first observe that, by (4.6), we have
$2^{j}\|\mathbf{1}_{A}\|_{2}\lesssim\|\Pi_{8\delta}^{2^{k}}\|_{2}.$ (4.9)
The radial decrease of the approximate identity implies that $\psi_{8\delta}\lesssim\psi_{8\delta}\ast\psi_{\delta}$, so
$\|\Pi^{2^{k}}_{8\delta}\|_{2}\lesssim\|\Pi^{2^{k}}\ast\psi_{8\delta}\ast\psi_{\delta}\|_{2}\leq\|\Pi^{2^{k}}\ast\psi_{\delta}\|_{2}\|\psi_{8\delta}\|_{1}=\|\Pi^{2^{k}}\|_{2}.$ (4.10)
Finally,
[…]
so $\delta^{\eta/5}\lesssim_{\eta}2^{j}\cdot\|\mathbf{1}_{A}\|_{1}\lesssim\Pi_{8\delta}^{2^{k}}(A)$ (by (4.6)). We have proved Claim 4.4(2) ∎
By Claim 4.4(2) and since $\operatorname{spt}\psi_{r}\subset B(r)$,
$\Pi^{2^{k}}\left([A]_{8r}\right)\geq\Pi_{8\delta}^{2^{k}}(A)\gtrsim\delta^{\eta/5},\text{ where }[A]_{8r}\text{ is the }8r-\text{neighbourhood of }A.$ (4.11)
@@
$(\nu*\sigma)([A]_{8\delta}+z)\gtrsim_{\eta}\delta^{\eta}/R^{2},\qquad z\in\mathbb{R}^{2}.$
So, recalling that $\eta=\epsilon/2$ with $\epsilon=\epsilon(\alpha,\beta,D)$, and taking $\delta=\delta(\eta,R)>0$ small enough,
$(\nu*\sigma)([A]_{8\delta}+z)\geq\delta^{\epsilon}.$
We also note that $\|\nu_{\delta}\|_{L^{2}}^{2}\leq\|\mu_{\delta}\|_{L^{2}}^{2}\leq\delta^{\alpha-2}$ by hypothesis. We may therefore apply Corollary 4.1 to the measures $\nu,\sigma$ and the set $E:=[A]_{8\delta}+z$, and conclude that
$|A|_{\delta}\geq\delta^{-\alpha-\epsilon}.$
So, by Claim 4.4(1), and again recalling $\epsilon=2\eta$,
$J(k)=\big\|\Pi^{2^{k}}\big\|_{2}\lesssim_{\eta}|A|_{\delta}^{-\frac{1}{2}}\delta^{-1-\eta/5}\leq\delta^{\frac{\alpha+\epsilon}{2}}\cdot\delta^{-1-\eta/5}\leq\delta^{\frac{\alpha+3\eta/2-2}{2}}.$
In particular $J(k)\leq\delta^{(\alpha+\eta-2)/2}$ for $\delta>0$ small enough. ∎
The next corollary follows by a straightforward iteration of Proposition 4.2:
@@
For all $D\geq 1$, $\mathfrak{d}>0$, $\beta\in[0,1)$, $t\in[0,2)$, there exist $k_{0}=k_{0}(D,\beta,t)\in\mathbb{N}$ and $\delta_{0}=\delta_{0}(D,\mathfrak{d},\beta,t)>0$ such that the following holds for all $\delta\in(0,\delta_{0}]$. Assume that $\sigma$ is a $(D,\beta)$-uniformly perfect probability measure supported on $\mathbb{P}$ with $\operatorname{diam}(\operatorname{spt}\sigma)\geq\mathfrak{d}$. Then,
$\|\sigma_{\delta}^{k}\|_{2}^{2}\leq\delta^{t-2},\qquad k\geq k_{0}.$
@@
Let $\alpha_{0}:=\tfrac{1}{2}\alpha$. The Frostman property easily implies $\|\sigma_{\delta}\|_{L^{2}}^{2}\lesssim_{D,\mathfrak{d},\beta}\delta^{\alpha-2}$, thus
$\|\sigma_{\delta}\|_{2}^{2}\leq\delta^{\alpha_{0}-2},\qquad\delta\in(0,\delta_{0}],$ (4.12)
for some $\delta_{0}=\delta_{0}(D,\mathfrak{d},\beta)>0$. We proceed to define a sequence of exponents $\{\alpha_{j}\}_{j=0}^{\infty}$, where $\alpha_{0}>0$ was already defined above. Given $j\geq 0$, we inductively define
@@
where $\eta_{j}:=\eta(\alpha_{j},\beta,D)>0$ is the constant provided by Proposition 4.2. Note that $\alpha_{j}\nearrow 2$ as $j\to\infty$, since $\eta(\alpha,\beta,D)>0$ stays bounded away from zero as $\alpha$ ranges on any fixed compact subset of $[0,2)$. In particular, given $t\in[0,2)$ as in the statement of the corollary, there exists $j_{0}=j_{0}(D,\beta,t)\in\mathbb{N}$ such that $\alpha_{j_{0}}\geq t$.
We make the following claim, to be proved by induction. Fix $j\geq 0$. Then, there exist $k^{j}=k^{j}(D,\beta,j)\in\mathbb{N}$ and $\delta^{j}=\delta^{j}(D,\mathfrak{d},\beta,j)>0$ such that
$\|\sigma_{\delta}^{k_{j}}\|_{2}^{2}\leq\delta^{\alpha_{j}-2},\qquad\delta\in(0,\delta^{j}].$ (4.13)
Applying this with $j:=j_{0}$, and setting $k_{0}:=k^{j_{0}}$, proves Corollary 4.5.
Let us then prove the inductive claim. The case $j=0$ follows from (4.12), with $k^{0}=1$. Let us then assume that the claim has already been established for some $j\geq 0$. Apply Proposition 4.2 with parameters $D,\beta$, $\alpha:=\alpha_{j}$, and radius
$R_{j}=2k^{j},$
where $k^{j}$ is the integer from the inductive hypothesis (4.13). The conclusion is the existence of $k_{0}=k_{0}(\alpha_{j},\beta,D)\in\mathbb{N}$, and $\delta_{0}:=\delta_{0}(\alpha_{j},\beta,\mathfrak{d},D,R_{j})>0$ such that the following holds for $\delta\in(0,\delta_{0}]$. If $\mu$ is a Borel probability measure satisfying
$\operatorname{spt}\mu\subset[-R_{j},R_{j}]^{2}\quad\text{and}\quad\|\mu_{\delta}\|_{L^{2}}^{2}\leq\delta^{\alpha_{j}-2},$ (4.14)
@@
$\|(\mu\ast\sigma)_{\delta}^{k_{0}}\|_{2}^{2}\leq\delta^{\alpha_{j}+\eta-2}=\delta^{\alpha_{j+1}-2}.$ (4.15)
But now by the inductive hypothesis (4.13), and since $\operatorname{spt}\sigma\subset[-2,2]^{2}$, the measure $\mu=\sigma^{k_{j}}$ satisfies (4.14) for $\delta\in(0,\delta^{j}]$. Therefore (4.15) yields
$\|\sigma_{\delta}^{k_{0}(k_{j}+1)}\|_{2}^{2}\leq\delta^{\alpha_{j+1}-2},\qquad\delta\in(0,\min\{\delta_{0},\delta^{j}\}].$
@@
Fix $R\geq 1$, and let $\{\varphi_{\delta}\}_{\delta>0}$ be an radially decreasing approximate identity with the property $|\widehat{\varphi_{\delta}}(\xi)|\gtrsim 1$ for $|\xi|\leq\delta$. Then, writing $\delta:=R^{-1}$, one has
$\|\hat{\sigma}\|_{L^{p}(B(R))}^{p}\lesssim\|\sigma^{k}_{\delta}\|_{2}^{2},\qquad p=2k\in 2\mathbb{N}.$ (4.16)
Write $t:=2-\epsilon$, where $\epsilon\in(0,1)$ is the parameter from the statement. Then, according to Corollary 4.5, we have $\|\sigma_{\delta}^{k}\|_{2}^{2}\leq\delta^{-\epsilon}$, provided $k\geq k_{0}(D,\beta,\epsilon)$, and $0<\delta\leq\delta_{0}(D,\mathfrak{d},\beta,\epsilon)$. This completes the proof in combination with (4.16). ∎